OpenStudy (anonymous):

What is the simplified form of the quantity 12 z-squared plus 25z plus 12 over the quantity 3 z-squared minus 2z minus 8 ? @sujuexob2st

OpenStudy (anonymous):

@C.A.M

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

Help on this one please: What is the simplified form of the quantity x squared plus 5x plus 6 over 15 x y squared all over the quantity 2x squared plus 7x plus 3 over 5 x squared y ?

OpenStudy (anonymous):

Give me a second

OpenStudy (anonymous):

can u use the equation tool bc its a bit confusing

OpenStudy (anonymous):

\[\frac{ \frac{ x ^{2}+5x+6 }{ 15xy ^{2} } }{ \frac{ 2x ^{2}+7x+3 }{ 5x ^{2} y} }\]

OpenStudy (anonymous):

\[\frac{ x^2+5x+6 }{ 15xy^2 } \times \frac{ 5x^2y }{ 2x^2+7x+3 }\]

OpenStudy (anonymous):

now factor and simplify

OpenStudy (anonymous):

can you show me please

OpenStudy (anonymous):

Do you know how to factor binomials and polynomials?

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

oh yes sorry about that

OpenStudy (anonymous):

So factor out x^2+5x+6 and 2x^2+7x+3 and find the common factor first

OpenStudy (anonymous):

Got it now thanks so much

OpenStudy (anonymous):

np